a spring makes 10 vibrations in 5 seconds. Find the period and frequency of the vibration.


1
Expert's answer
2022-01-31T07:33:52-0500

Period:

"T=1\ vibration×\frac{5\ s}{10\ vibrations}=0.5\ s"


Frequency:

"f=\frac{1}{T}=\frac{1}{0.5\ s}=2\ Hz"